Hotel rooms in Smalltown go for $100, and

1,000 rooms are rented on a typical day.

a. To raise revenue, the mayor decides to charge

hotels a tax of $10 per rented room. After

the tax is imposed, the going rate for hotel

rooms rises to $108, and the number of rooms

rented falls to 900. Calculate the amount of

revenue this tax raises for Smalltown and the

deadweight loss of the tax. (Hint: The area of

a triangle is 1⁄2 3 base 3 height.)

b. The mayor now doubles the tax to $20. The

price rises to $116, and the number of rooms

rented falls to 800. Calculate tax revenue

and deadweight loss with this larger tax. Do

they double, more than double, or less than

double? Explain.


1
Expert's answer
2022-02-02T09:53:46-0500

The tax revenue is calculated by :


Tax revenue = Tax Imposed "\times" Quantity sold

Tax revenue = $10 "\times" 900

= $ 9000


The deadweight loss is calculated as:


Deadweight loss = "\frac{1}{2}" "\times" Tax imposed "\times" change in quantity sold


Deadweight loss = "\frac{1}{2}" "\times" $10 "\times" (1000 - 900)

= $ 500


Hence,

Tax revenue = $ 9000

Deadweight loss = $ 500


PART(B):


Tax revenue = Tax imposed "\times" quantity sold

= $ 20 "\times" 800

= $ 16000


Deadweight loss = "\frac{1}{2}" "\times" Tax imposed "\times" change in quantity sold


= "\frac{1}{2}" "\times" $ 20 "\times" (1000 - 800)

= $ 2000


Therefore;


When a tax rate of $20 is imposed, it leads to the following:


Tax revenue rises to $ 16000

Deadweight loss becomes $ 2000

Tax revenue is less than double as it rises from $ 9000 to $ 16000

Deadweight loss is more than double as it rises from $ 500 to $ 2000.